# Matérn Kernels for Tunable Implicit Surface Reconstruction

### v2

@@ 4 Experiments and Results
We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (Chang et al., 2015) and the Surface Reconstruction Benchmark (SRB; Berger et al. (2013)) in Section 4.1. Moreover, Section 4.2 demonstrates Matérn kernels’ ability to reconstruct high-frequency textures on the Google Scanned Objects (GSO; Downs et al. (2022)) dataset and Objaverse (Deitke et al., 2023). Leveraging Neural Kernel Fields (Williams et al., 2022), we present and evaluate learnable Matérn kernels in Section 4.3.
@@ 4.1 Surface Reconstruction on ShapeNet and SRB
ShapeNet. We compare Matérn kernels in a sparse setting against the arc-cosine kernel on ShapeNet (using train/val/test split provided by Mescheder et al. (2019)). To do so, we randomly sample $m=1,000$ surface points with corresponding normals for each shape. We implemented Matérn kernels in PyTorch and took the official CUDA implementation of the arc-cosine kernel from NS, eventually integrated into a unified framework to ensure fair comparison. Notably, we did not use Nyström sampling as in NS and employed PyTorch’s direct Cholesky solver to solve the KRR problem in Eq. (1) instead of an iterative conjugate gradient solver. Moreover, similar to NS, we approximate the gradient part in Eq. (1) with finite differences. See Appendix F for details.
@@
[…]
Quantitative results in terms of F-Score (with a threshold of 0.01) and Chamfer distance (CD; always reported $\times$ 103) can be found in Table 1. Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (Kazhdan & Hoppe, 2013). In addition, our experiments show that the bandwidth parameter, $h$, can be used to tune the surface reconstruction (lower the error), while not being too sensitive in practice. Matérn 5/2 and the Gaussian kernel (for $\nu\rightarrow\infty$) perform significantly worse, being simply too smooth. Figure 2 and Appendix F shows qualitative results.
@@
SRB. Next, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark which consists of five complex shapes simulated from incomplete and noisy range scans with up to 100,000 points. For this, we implemented a highly scalable version of Matérn kernels, closely following the NS framework to solve the KRR problem in Eq. (1) using Nyström sampling (with 15,000 samples) and the FALKON (Rudi et al., 2017) solver. We compare against NS and SPSR, as well as kernel-free methods, Fourier Feature Networks (FFNs; Tancik et al. (2020)), Siren, SAP (Peng et al., 2021), DiGS (Ben-Shabat et al., 2022), VisCo (Pumarola et al., 2022), and OG-INR (Koneputugodage et al., 2023). All methods are optimization-based (FFN and Siren are overfitted to a single shape) and use surface normals. For Matérn kernels, we did a modest parameter sweep over $h\in\{0.5,1,2\}$ and took the reconstruction with the lowest Chamfer distance. Runtime is measured on a single NVIDIA V100.
@@ 4.2 Texture Reconstruction on GSO and Objaverse
We also show qualitative results on the challenging Objaverse (Deitke et al., 2023) dataset in Figure 7. We chose Objaverse as it includes extremely high-resolution and complex textures. Matérn 3/2 reconstructs high-frequency texture details with great precision, overall yielding visually more pleasant reconstructions than the arc-cosine kernel. Reconstructed textures have fewer artifacts and are generally sharper, demonstrating that Matérn kernels can overcome spectral bias much better than the arc-cosine kernel. Note that this perfectly confirms our theoretical analysis presented in Section 3.2.
@@ 4.3 Data-dependent Matérn Kernels
We leverage the Neural Kernel Field (NKF) framework introduced by Williams et al. (2022) to make Matérn kernels learnable. Specifically, NKF proposed to feed points $x,y\in\mathbb{R}^{d}$ through an input-conditioned neural network $\varphi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{q}$ before evaluating a kernel function, $k^{\varphi}(x,y;\mathcal{D})=k\left([x,\varphi(x;\mathcal{D})],[y,\varphi(y;\mathcal{D})]\right)$. We re-implemented the NKF framework since the authors did not provide code. We set $q=32$ for our experiments. Please see Williams et al. (2022) for details.
Sparse surface reconstruction and extreme generalization. We compare learned Matérn kernels against the original NKF framework (with the arc-cosine kernel) and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet. Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels. Additionally, we evaluate Matérn kernels’ out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories. Table 4 reports the results, quantified using intersection-over-union (IoU), Chamfer distance, and normal consistency (NC). Learned Matérn 1/2 and 3/2 outperform the arc-cosine kernel while being significantly faster to train. Training NKF on the entire ShapeNet dataset (consisting of approx. 30,000 shapes) takes about six days—almost twice as long as for Matérn kernels, which require about three days (measured on a single NVIDIA A100). Additionally, although Matérn 1/2 is not able to surpass NKSR, it comes very close (95.3 vs. 95.6 in IoU) while having a shorter training time. Regarding out-of-category generalization, we observe that Matérn 1/2 performs best, followed by Matérn 3/2, NKF, and NKSR. Please see Figure 6 (right) and Appendix I for qualitative results.
@@
[…]
@@ Appendix A On the Stationarity of Matérn Kernels
As noted in Section 3.1 of the main paper, Matérn kernels are stationary: they only depend on the difference $x-y$ between two points $x,y\in\Omega$, hence being translation invariant. This is because $k(x-y)=k(x+t-(y+t))$ for $t\in\mathbb{R}^{d}$ and all stationary kernels, $k$. In contrast, the arc-cosine kernel is not translation invariant as its value depends on the absolute positions, $x,y$, see Eq. (3) of the main paper (i.e., the arc-cosine kernel can not be written as a function of $x-y$). In fact, the arc-cosine kernel is non-stationary. In the following and with the help of Figure 8, we will further discuss why stationary kernels, such as Matérn kernels, might be a better choice for kernel-based implicit surface reconstruction than non-stationary kernels, such as the arc-cosine kernel.
@@ Appendix D Proof of Theorem 3
because $k$ is real and for even spectral densities. On the other hand, Rahimi & Recht (2007) showed
@@
where ()(*)( ∗ ) follows because $b$ is uniformly distributed in $[0,2\pi]$ (hence, the inner expectation w.r.t. $b$ becomes zero).
@@
with bottom-layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$, biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$, and top-layer weights $v=(v_{1},v_{2},\dots,v_{m})\in\mathbb{R}^{m}$. Assume that $W$ and $b$ are fixed (that is, we only allow the second layer to be trained), and initialized according to some distribution. Choose
@@
${=}k_{\nu}(x,y).$
Here, ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see, e.g., Eq. (10) in Kanagawa et al. (2018). ∎
@@
The NTK of a Siren can be associated with any stationary kernel if the bottom-layer weights in Eq. (13) are initialized according to the corresponding kernel’s spectral density.
@@ Appendix E Proof of Corollary 5
Theorem 8 (Saitoh & Sawano (2016), Theorem 2.17).
@@
(1) Follows from Theorem 4 by noting that the NTK of a fully-connected $(L=2$)-layer ReLU network is the first-order arc-cosine kernel, see, e.g.Cho & Saul (2009); Williams et al. (2021). For (2), we know that $H_{\nu}\subset H_{\text{AC}}$ and $H_{\text{AC}}\subset H_{\nu}$, so, based on Theorem 8, $k:=\gamma^{2}k_{\text{AC}}-k_{\nu}$ and $\bar{k}:=\bar{\gamma}^{2}k_{\nu}-k_{\text{AC}}$ are both valid kernels for $\gamma,\bar{\gamma}>0$. Re-arranging both equations, we have
$k_{\text{AC}}=ak_{\nu}+b,\quad\text{where}\quad a=1/\gamma^{2},b=k/\gamma^{2}$
and, equivalently,
$k_{\nu}=\bar{a}k_{\text{AC}}+\bar{b},\quad\text{where}\quad\bar{a}=1/\bar{\gamma}^{2},\bar{b}=\bar{k}/\bar{\gamma}^{2}.$
This shows that the Laplace and arc-cosine kernel are indeed related by an affine transformation when restricted to $\mathbb{S}^{d-1}$, concluding the proof. ∎
@@ F.1 Implementation Details
Following common practice (see, e.g., Williams et al. (2022)), we use finite differences to approximate the gradient part of the KRR problem in Eq. (1) of the main paper. Specifically, denote $\mathcal{X}^{\prime}:=\mathcal{X}^{+}\cup\mathcal{X}^{-}$, where $\mathcal{X}^{+}:=\{x_{1}+{\epsilon}n_{1},x_{2}+{\epsilon}n_{2},\dots,x_{m}+{\epsilon}n_{m}\}$ and $\mathcal{X}^{-}:=\{x_{1}-{\epsilon}n_{1},x_{2}-{\epsilon}n_{2},\dots,x_{m}-{\epsilon}n_{m}\}$ for a fixed ${\epsilon}>0$. The Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) tells us that the solution to Eq. (1) is of the form
@@ Appendix G Further Details on the Surface Reconstruction Benchmark
This section details the experimental setting that has been used to run Neural Splines (NS; Williams et al. (2021)) on SRB in Section 4.1 of the main paper. Moreover, we present per-object metrics and additional qualitative results.
@@ G.1 Experimental Setting
Despite our best efforts, we were unable to reproduce the results for the NS kernel reported in the original paper. We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines), and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance. In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; Williams et al. (2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively. In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
To measure runtime, we again used the same setting as in the original NS paper: 15,000 Nyström samples, and $\lambda=10^{-11}$. Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters. Runtime was measured on a single NVIDIA V100, similar to Williams et al. (2021).
@@ G.2 Additional Results
Figure 10: Additional qualitative results on SRB for non-learnable kernels. We compare Neural Splines (NS; Williams et al. (2021)) to Matérn 1/2 and 3/2.
@@ H.1 Implementation Details
To extract a textured surface mesh from the predicted four-dimensional volume, we use a two-stage process. First, we apply Marching Cubes (Lorensen & Cline, 1987) to $f^{\prime}_{1}$, and then trilinearily interpolate RGB colors at the previously predicted surface points.
@@ H.2 Evaluation Protocol
Our quantitative evaluation on the Google Scanned Objects (GSO; Downs et al. (2022)) dataset closely follows Mitchel et al. (2024), in which authors compute image-based evaluation metrics on the texture atlas to quantify reconstruction ability. Specifically, we first sample $m$ surface points along with corresponding normals and per-point RGB colors from shapes in GSO. Next, we solve the KRR problem in Eq. (1) as always. Then, instead of predicting RGB colors on the estimated geometry (which generally differs between various kernel functions), we predict (i.e., trilinearily interpolate) colors on the surface points of the original mesh. This allows us to leverage the existing UV parametrization of the original mesh to generate a texture atlas with colors predicted from KRR. Once the texture atlas at hand, we compare it to the original texture atlas using PSNR and LPIPS.

### v3

@@ 4 Experiments and Results
We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (chang2015) and the Surface Reconstruction Benchmark (SRB; berger2013) in Section 4.1. Moreover, Section 4.2 demonstrates Matérn kernels’ ability to reconstruct high-frequency textures on the Google Scanned Objects (GSO; downs2022) dataset and Objaverse (deitke2023). Leveraging Neural Kernel Fields (williams2022), we present and evaluate learnable Matérn kernels in Section 4.3.
@@ 4.1 Surface Reconstruction on ShapeNet and SRB
ShapeNet. We compare Matérn kernels in a sparse setting against the arc-cosine kernel on ShapeNet (using train/val/test split provided by mescheder2019). To do so, we randomly sample $m=1,000$ surface points with corresponding normals for each shape. We implemented Matérn kernels in PyTorch and took the official CUDA implementation of the arc-cosine kernel from NS, eventually integrated into a unified framework to ensure fair comparison. Notably, we did not use Nyström sampling as in NS and employed PyTorch’s direct Cholesky solver to solve the KRR problem in Eq. (1) instead of an iterative conjugate gradient solver. Moreover, similar to NS, we approximate the gradient part in Eq. (1) with finite differences. See Appendix F for details.
@@
Table 1: Results on ShapeNet (left) and SRB (right) for non-learnable kernels. Arc-cosine kernel and Matérn $\nu\in\{\text{1/2},\text{3/2},\text{5/2},\infty\}$ are implemented in a unified framework and based on the exact same parameters. ”NS*” denotes the best result we could achieve by running the official implementation of NS, see Appendix G. Bold marks best result, underline second best.
Quantitative results in terms of F-Score (with a threshold of 0.01) and Chamfer distance (CD; always reported $\times$ 103) can be found in Table 1. Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (kazhdan2013). In addition, our experiments show that the bandwidth parameter, $h$, can be used to tune the surface reconstruction (lower the error), while not being too sensitive in practice. Matérn 5/2 and the Gaussian kernel (for $\nu\rightarrow\infty$) perform significantly worse, being simply too smooth. Figure 2 and Appendix F shows qualitative results.
@@
SRB. Next, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark which consists of five complex shapes simulated from incomplete and noisy range scans with up to 100,000 points. For this, we implemented a highly scalable version of Matérn kernels, closely following the NS framework to solve the KRR problem in Eq. (1) using Nyström sampling (with 15,000 samples) and the FALKON (rudi2017) solver. We compare against NS and SPSR, as well as kernel-free methods, Fourier Feature Networks (FFNs; tancik2020), Siren, SAP (peng2021), DiGS (shabat2022), VisCo (pumarola2022), and OG-INR (koneputugodage2023). All methods are optimization-based (FFN and Siren are overfitted to a single shape) and use surface normals. For Matérn kernels, we did a modest parameter sweep over $h\in\{0.5,1,2\}$ and took the reconstruction with the lowest Chamfer distance. Runtime is measured on a single NVIDIA V100.
@@ 4.2 Texture Reconstruction on GSO and Objaverse
We also show qualitative results on the challenging Objaverse (deitke2023) dataset in Figure 7. We chose Objaverse as it includes extremely high-resolution and complex textures. Matérn 3/2 reconstructs high-frequency texture details with great precision, overall yielding visually more pleasant reconstructions than the arc-cosine kernel. Reconstructed textures have fewer artifacts and are generally sharper, demonstrating that Matérn kernels can overcome spectral bias much better than the arc-cosine kernel. Note that this perfectly confirms our theoretical analysis presented in Section 3.2.
@@ 4.3 Data-dependent Matérn Kernels
We leverage the Neural Kernel Field (NKF) framework introduced by williams2022 to make Matérn kernels learnable. Specifically, NKF proposed to feed points $x,y\in\mathbb{R}^{d}$ through an input-conditioned neural network $\varphi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{q}$ before evaluating a kernel function, $k^{\varphi}(x,y;\mathcal{D})=k\left([x,\varphi(x;\mathcal{D})],[y,\varphi(y;\mathcal{D})]\right)$. We re-implemented the NKF framework since the authors did not provide code. We set $q=32$ for our experiments. Please see williams2022 for details.
Sparse surface reconstruction and extreme generalization. We compare learned Matérn kernels against the original NKF framework (with the arc-cosine kernel) and NKSR (huang2023) in a sparse setting on ShapeNet. Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels. Additionally, we evaluate Matérn kernels’ out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories. Table 4 reports the results, quantified using intersection-over-union (IoU), Chamfer distance, and normal consistency (NC). Learned Matérn 1/2 and 3/2 outperform the arc-cosine kernel while being significantly faster to train. Training NKF on the entire ShapeNet dataset (consisting of approx. 30,000 shapes) takes about six days—almost twice as long as for Matérn kernels, which require about three days (measured on a single NVIDIA A100). Additionally, although Matérn 1/2 is not able to surpass NKSR, it comes very close (95.3 vs. 95.6 in IoU) while having a shorter training time. Regarding out-of-category generalization, we observe that Matérn 1/2 performs best, followed by Matérn 3/2, NKF, and NKSR. Please see Figure 6 (right) and Appendix I for qualitative results.
@@
Table 4: Results on ShapeNet for learned kernels. ”NKF*” denotes a re-implemented variant of NKF. NKF* and Matérn $\nu\in\{\text{1/2},\text{3/2},\text{5/2},\infty\}$ are based on the same framework and parameters; they differ only in the employed kernel. Runtime is measured on a single NVIDIA A100 with a batch size of one to ensure fair comparison. Bold marks best result, underline second best.
@@ Appendix A On the Stationarity of Matérn Kernels
As noted in Section 3.1 of the main paper, Matérn kernels are stationary: they only depend on the difference xyx-y between two points $x,y\in\Omega$, hence being translation invariant. This is because k(xy)=k(x+t(y+t))k(x-y)=k(x+t-(y+t)) for $t\in\mathbb{R}^{d}$ and all stationary kernels, $k$. In contrast, the arc-cosine kernel is not translation invariant as its value depends on the absolute positions, $x,y$, see Eq. (3) of the main paper (i.e., the arc-cosine kernel can not be written as a function of xyx-y). In fact, the arc-cosine kernel is non-stationary. In the following and with the help of Figure 8, we will further discuss why stationary kernels, such as Matérn kernels, might be a better choice for kernel-based implicit surface reconstruction than non-stationary kernels, such as the arc-cosine kernel.
@@ Appendix D Proof of Theorem 3
because $k$ is real and for even spectral densities. On the other hand, rahimi2007 showed
@@
where ()(*) follows because $b$ is uniformly distributed in $0,2\pi]$ (hence, the inner expectation w.r.t. $b$ becomes zero).
@@
with bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$, biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$, and top-layer weights $v=(v_{1},v_{2},\dots,v_{m})\in\mathbb{R}^{m}$. Assume that $W$ and $b$ are fixed (that is, we only allow the second layer to be trained), and initialized according to some distribution. Choose
@@
$\hskip-10.27777pt\overset{(**)}{=}k_{\nu}(x,y).$
Here, ()(*) follows because $b$ is uniformly distributed, and ()(**) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see, e.g., Eq. (10) in kanagawa2018. ∎
@@
The NTK of a Siren can be associated with any stationary kernel if the bottom layer weights in Eq. (13) are initialized according to the corresponding kernel’s spectral density.
@@ Appendix E Proof of Corollary 5
Theorem 8 (saitoh2016, Theorem 2.17).
@@
The first part follows from Theorem 4 by observing that the NTK of a fully-connected $(L=2$)-layer ReLU network (with all parameters being trained) is
$k_{\widetilde{\text{AC}}}(x,y)=\underbrace{k_{\text{AC}}(x,y)}_{\text{first-order}}+x^{\top}y\underbrace{\frac{1}{\pi}(\pi-\theta)}_{\text{zeroth-order}}$ (14)
see, e.g.chen2021. The second part follows from Theorem 4, noting that $H_{\widetilde{\text{AC}}}\subset H_{\text{1/2}}$ when restricted to $\mathbb{S}^{d-1}$; so, based on Theorem 8, $k:=\gamma^{2}k_{\text{1/2}}-k_{\widetilde{\text{AC}}}$ is a valid kernel for $\gamma>0$. Re-arranging the equation, we have
$k_{\text{1/2}}=ak_{\widetilde{\text{AC}}}+b,\quad\text{where}\quad a=1/\gamma^{2},b=k/\gamma^{2}.$ (15)
Now, denote by $k_{\text{AC}}^{(0)}(x,y)$ the zeroth-order arc-cosine kernel as defined in Eq. (14), and further set $\hat{k}_{\text{AC}}^{(0)}(x,y)=x^{\top}yk_{\text{AC}}^{(0)}(x,y)$. The latter is a product between two kernel functions, hence a valid, i.e., positive definite, kernel itself. Plugin Eq. (14) into Eq. (15) yields
$\begin{split}k_{\text{1/2}}&=a(k_{\text{AC}}+\hat{k}_{\text{AC}}^{(0)})+b\\ &=ak_{\text{AC}}+c\end{split}\quad\text{with}\quad c=a\hat{k}_{\text{AC}}^{(0)}+b,$ (16)
where $c$ is again a valid kernel function since it is a sum of two kernels. This shows that the Laplace and first-order arc-cosine kernel are indeed related by a constant scaling factor $a$ and the addition of a kernel function $c$ when restricted to $\mathbb{S}^{d-1}$, concluding the proof. ∎
In other words, Corollary 5 states that, when restricted to the sphere, the Laplace kernel is a linear combination of two kernel functions: the first-order arc-cosine kernel and another, valid kernel function (that accounts for the residuals). This holds also in the other direction.
@@ F.1 Implementation Details
Following common practice (see, e.g., williams2022), we use finite differences to approximate the gradient part of the KRR problem in Eq. (1) of the main paper. Specifically, denote $\mathcal{X}^{\prime}:=\mathcal{X}^{+}\cup\mathcal{X}^{-}$, where $\mathcal{X}^{+}:=\{x_{1}+{\epsilon}n_{1},x_{2}+{\epsilon}n_{2},\dots,x_{m}+{\epsilon}n_{m}\}$ and $\mathcal{X}^{-}:=\{x_{1}-{\epsilon}n_{1},x_{2}-{\epsilon}n_{2},\dots,x_{m}-{\epsilon}n_{m}\}$ for a fixed ${\epsilon}>0$. The Representer Theorem (kimeldorf1970; schoelkopf2001) tells us that the solution to Eq. (1) is of the form
@@ Appendix G Further Details on the Surface Reconstruction Benchmark
This section details the experimental setting that has been used to run Neural Splines (NS; williams2021) on SRB in Section 4.1 of the main paper. Moreover, we present per-object metrics and additional qualitative results.
@@ G.1 Experimental Setting
Despite our best efforts, we were unable to reproduce the results for the NS kernel reported in the original paper. We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines), and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance. In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; williams2021 reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively. In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
To measure runtime, we again used the same setting as in the original NS paper: 15,000 Nyström samples, and $\lambda=10^{-11}$. Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters. Runtime was measured on a single NVIDIA V100, similar to williams2021.
@@ G.2 Additional Results
Figure 10: Additional qualitative results on SRB for non-learnable kernels. We compare Neural Splines (NS; williams2021) to Matérn 1/2 and 3/2.
@@ H.1 Implementation Details
To extract a textured surface mesh from the predicted four-dimensional volume, we use a two-stage process. First, we apply Marching Cubes (lorensen1987) to $f^{\prime}_{1}$, and then trilinearily interpolate RGB colors at the previously predicted surface points.
@@ H.2 Evaluation Protocol
Our quantitative evaluation on the Google Scanned Objects (GSO; downs2022) dataset closely follows mitchel2024, in which authors compute image-based evaluation metrics on the texture atlas to quantify reconstruction ability. Specifically, we first sample $m$ surface points along with corresponding normals and per-point RGB colors from shapes in GSO. Next, we solve the KRR problem in Eq. (1) as always. Then, instead of predicting RGB colors on the estimated geometry (which generally differs between various kernel functions), we predict (i.e., trilinearily interpolate) colors on the surface points of the original mesh. This allows us to leverage the existing UV parametrization of the original mesh to generate a texture atlas with colors predicted from KRR. Once the texture atlas at hand, we compare it to the original texture atlas using PSNR and LPIPS.